# Weak Cosmic Censorship and Weak Gravity Conjectures in CFT Thermodynamics

### v1

@@ 3 The Model
Reissner-Nordström AdS black holes, which are enveloped by PFDM, carry an electrical charge and are affected by a theoretical type of matter that constitutes the majority of the universe’s mass. These black holes exhibit intriguing thermodynamic behaviors and phase transition characteristics, which can be explored by broadening the phase space of the system. Additionally, some studies have investigated the WCCC in the context of these black holes. This conjecture posits that a black hole’s singularity remains hidden from any external observer. The action under consideration pertains to a gravitational theory that interacts minimally with a gauge field amidst the presence of PFDM[105, 106, 107, 108, 109],
$S=\int d^{4}x\sqrt{-g}\bigg{[}\frac{1}{16\pi G}R+\frac{1}{4}F^{\mu\nu}F_{\mu\nu}-\frac{\Lambda}{8\pi G}+\mathcal{L}_{DM}\bigg{]},$ (1)
@@
$\begin{split}R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=&-8\pi G(T^{M}_{\mu\nu}+T^{DM}_{\mu\nu})\equiv-8\pi GT_{\mu\nu},\end{split}$ (2)
$\begin{split}&F^{\mu\nu}_{\nu}=0,\\ &F^{\mu\nu;\alpha}+F^{\nu\alpha;\mu}+F^{\alpha\mu;\nu}=0.\end{split}$ (3)
The energy-momentum tensor for ordinary matter is symbolized by $T^{M}_{\mu\nu}$, and the energy-momentum tensor for PFDM is represented by $T^{DM}_{\mu\nu}$. These tensors are detailed in[105, 106, 107, 108, 109]. Consequently, the equations can be expressed as follows,
$\begin{split}&T^{\mu}_{\nu}=g^{\mu\nu}T_{\mu\nu},\\ &T^{t}_{t}=-\rho,\hskip 14.22636ptT^{r}_{r}=T^{\theta}_{\theta}=T^{\phi}_{\phi}=P.\end{split}$ (4)
In the framework of the RN AdS metric influenced by PFDM, it is posited that the components of the energy-momentum tensor satisfy $T^{r}_{r}=T^{\theta}_{\theta}=T^{\phi}_{\phi}=T^{t}_{t}(1-\delta)$, with $\delta$ being a constant. This assumption is substantiated by [105, 106, 107, 108, 109]. The formulation of the RN AdS metric within the PFDM context is delineated as follows,
$\begin{split}ds^{2}=-f(r)dt^{2}+f(r)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),\end{split}$ (5)
@@
$f(r)=1-\frac{2MG}{r}+\frac{GQ^{2}}{r^{2}}-\frac{\Lambda}{3}r^{2}+\frac{\gamma}{r}\ln(\frac{r}{\gamma}).$ (6)
In this analysis, the parameters $M$, $Q$, and $\gamma$ signify the mass, charge and the influence of (PFDM) on the black hole, respectively. Also, $\Lambda=-3/\ell^{2}$ is the cosmological constant with $\ell$ being the AdS radius. In the absence of PFDM ($\gamma=0$), the space-time metric simplifies to that of a standard RN AdS black hole. The parameter $\gamma$ can assume both positive and negative values, reflecting different PFDM scenarios. This study contemplates the implications of both positive and negative $\gamma$ on the black hole’s characteristics.
@@ 4 RPS thermodynamics
In the RPS formalism, the traditional variables of pressure $P$ and volume $V$ are eschewed due to the altered connotations they acquire within the holographic framework. Instead, two pivotal quantities, $C$ and $\mu$, assume significant roles within the realm of CFT. The central charge $C$ is instrumental in ascertaining the microscopic degrees of freedom inherent in CFT, while its conjugate $\mu$ is regarded as a chemical potential, as referenced in literature[20]. Within the domain of extended phase space thermodynamics, the exploration of black hole thermodynamics necessitates the employment of a dynamic cosmological constant. This adjustment invariably leads to a transformation in the associated dynamical equations. To circumvent this modification, the thermodynamics of a constrained phase space is proposed, wherein Newton’s constant is treated as a variable, supplanting the cosmological constant. This substitution preserves the original form of the field equations. Moreover, a correlation is established between the central charge and the cosmological and Newton’s constants, expressed as $C=\frac{\ell^{2}}{G}$, where $\ell^{2}=-\frac{3}{\Lambda}$ [20]. The first law of thermodynamics, as applied to a RN-AdS black hole surrounded by PFDM within the RPS framework, is articulated as follows,
$dM=TdS+\hat{\phi}d\hat{Q}+\mu dC+\Pi d\gamma.$ (7)
The observation aligns with an Euler-like relation, which is a fundamental principle in thermodynamics and fluid dynamics, often used to describe the relationship between various state functions. In the context of the RPS formalism, such relations are crucial for understanding the thermodynamic properties of black holes and their behavior in different spacetime geometries. By incorporating this Euler-like relation into the RPS framework, researchers can further elucidate the intricate dynamics of black holes within the holographic paradigm, offering deeper insights into the nature of spacetime and the fundamental forces that govern it,
$M=TS+\hat{\phi}\hat{Q}+\mu C+\Pi\gamma,$ (8)
in the context of our black hole, the parameters $M$, $T$, and $S$ represent the mass, temperature, and entropy, respectively. Moreover, these parameters can generally be characterized by corresponding quantities in the dual CFT,
$\hat{Q}=\frac{Q\ell}{\sqrt{G}}\hskip 28.45274pt\hskip 28.45274pt\hat{\phi}=\frac{\phi\sqrt{G}}{\ell}.$ (9)
@@
$C=\frac{\ell^{2}}{G},\qquad\hat{Q}=Q\sqrt{C},\qquad r_{+}=\ell\sqrt{\frac{S}{\pi C}}.$ (10)
Setting $(f(r_{+})=0)$ and utilizing equations (6), (7), (9), and (10), we can derive some variables such as M, T, $\hat{\phi}$, $\Pi$, and $\mu$ of the black hole as follows,
$M=\frac{\pi CS+\pi^{2}\hat{Q}^{2}+S^{2}}{2\pi^{3/2}\ell\sqrt{CS}}+\frac{C\gamma}{2\ell^{2}}\ln\bigg{(}\sqrt{\frac{S}{\pi C}}\frac{\ell}{\gamma}\bigg{)},$ (11)
$T=\bigg{(}\frac{\partial M}{\partial S}\bigg{)}_{\hat{Q},C,\gamma}=\frac{\pi CS-\pi^{2}\hat{Q}^{2}+3S^{2}}{4\pi^{3/2}\ell S\sqrt{CS}}+\frac{C\gamma}{4\ell^{2}S}$ (12)
$\hat{\phi}=\bigg{(}\frac{\partial M}{\partial\hat{Q}}\bigg{)}_{S,C,\gamma}=\sqrt{\frac{\pi}{SC}}\frac{\hat{Q}}{\ell},$ (13)
$\Pi=\bigg{(}\frac{\partial M}{\partial\gamma}\bigg{)}_{S,\hat{Q},C}=\frac{C}{2\ell^{2}}\bigg{[}-1+\ln\bigg{(}\sqrt{\frac{S}{\pi C}}\frac{\ell}{\gamma}\bigg{)}\bigg{]},$ (14)
$\mu=\bigg{(}\frac{\partial M}{\partial\gamma}\bigg{)}_{S,\hat{Q},\gamma}=\frac{\pi CS-\pi^{2}\hat{Q}^{2}-S^{2}}{4\pi^{3/2}C\ell\sqrt{CS}}+\frac{\gamma}{4\ell^{2}}\bigg{[}-1+2\ln\bigg{(}\sqrt{\frac{S}{\pi C}}\frac{\ell}{\gamma}\bigg{)}\bigg{]}.$ (15)
@@
$\bigg{(}\frac{\partial T}{\partial S}\bigg{)}_{\hat{Q},C,\gamma}=0,\qquad\bigg{(}\frac{\partial^{2}T}{\partial S^{2}}\bigg{)}_{\hat{Q},C,\gamma}=0.$ (16)
Using Eqs. (12) and (16) for the critical parameters, we have,
$Y:=C\ell^{\frac{4}{3}}\big{(}81\gamma^{2}-\ell^{2}+9\ell\gamma\sqrt{81\gamma^{2}-2\ell^{2}}\big{)}^{\frac{1}{3}},\qquad S_{c}=\frac{\pi(C\ell^{2}+Y)^{2}}{18\ell^{2}Y},$ (17)
$\hat{Q}_{c}^{2}=-\frac{\left(C\ell^{2}+Y\right)^{2}\left(C^{2}\ell^{4}-4C\ell^{2}Y+Y^{2}\right)}{324\ell^{4}Y^{2}}+\frac{\sqrt{2}\gamma C^{3/2}\left(C\ell^{2}+Y\right)}{9\sqrt{Y}\ell^{2}}.$ (18)
@@
$T_{c}=\frac{\left(C\ell^{2}+Y\right)\left(-9\sqrt{2}C^{3/2}\gamma\ell^{2}Y^{3/2}+C^{3}\ell^{6}+6C^{2}\ell^{4}Y+6C\ell^{2}Y^{2}+Y^{3}\right)}{3\pi\sqrt{2CY}\ell^{2}\left(C\ell^{2}+Y\right)^{3}}+\frac{9\gamma CY}{2\pi\left(C\ell^{2}+Y\right)^{2}},$ (19)
$M_{c}=\frac{\sqrt{C}\left(C\ell^{2}+Y\right)\left(2C\ell^{2}+3\gamma\sqrt{2CY}+2Y\right)}{9\ell^{2}\sqrt{2Y}\left(C\ell^{2}+Y\right)}+\frac{\gamma C}{2\ell^{2}}\ln\left(\frac{C\ell^{2}+Y}{\gamma\sqrt{18CY}}\right).$ (20)
@@
$F=M-TS=\frac{\pi CS+3\pi^{2}\hat{Q}^{2}-S^{2}}{4\pi^{3/2}\ell\sqrt{CS}}+\frac{\gamma C}{4\ell^{2}}\bigg{[}-1+2\ln\bigg{(}\frac{\ell}{\gamma}\sqrt{\frac{S}{\pi C}}\bigg{)}\bigg{]}.$ (21)
Figures 1 and 2 depict the evolution of free energy as a function of temperature for various free parameter values of the RN AdS black hole, analyzed within the framework of PFDM in RPS. When $\hat{Q}_{c}>\hat{Q}=0.5$, a distinctive swallowtail pattern emerges, indicative of a first-order phase transition. Also, for ($\hat{Q}_{c}=\hat{Q}$ ), we have a second-order phase transition. To enhance clarity, we have separated the figures. The specific free parameters pertinent to each figure are duly annotated. The separation of the figures isolates and emphasizes each parameter set’s unique characteristics, facilitating a more detailed examination of the system’s behavior under varying conditions. It is essential to recognize that the free parameters play a crucial role in shaping the physical properties of the black hole.
Figure 1: With the $\gamma=2,C=\ell=1$ and $\hat{Q}_{c}=0.93$ (a) $\hat{Q}_{c}>\hat{Q}=0.5$ (b) $\hat{Q}_{c}=\hat{Q}=0.93$ (c) $\hat{Q}_{c}<\hat{Q}=1.5$
Figure 2: With the $\gamma=1.5,C=\ell=1$ and $\hat{Q}_{c}=0.79$ (a) $\hat{Q}_{c}>\hat{Q}=0.5$ (b) $\hat{Q}_{c}=\hat{Q}=0.79$ (c) $\hat{Q}_{c}<\hat{Q}=1.5$
Figure 3: spinodal curve ($T-S$) for charged black hole in presence perfect fluid. We have set $C=\ell=1$ and (a) $\gamma=2$ and $\hat{Q}_{c}=0.93$, (b) $\gamma=1.5$ and $\hat{Q}_{c}=0.79$
Figure 3 presents the Hawking temperature as a function of entropy for the AdS (RN) black hole amidst the presence of (PFDM) in the (RPS). We have set the constants $C$ and $\ell$ to 1. For Figure 3a, the parameters are set to $\gamma=2$ and $\hat{Q}_{c}$= 0.93 , while for Figure 3b, $\gamma=1.5$ and $\hat{Q}_{c}$= 0.79 . Each subplot within the figure reveals certain zero points for varying free parameters within the range $10<S<$ 1. Upon examining the entropy alongside the black hole’s free parameter values, it is observed that for small values of $S$, the structural behavior of the black hole is markedly unique. Conversely, as the value of $S$ increases, the figures tend to converge, indicating a unification in behavior for larger entropy values. This convergence suggests that the influence of the free parameters diminishes as the entropy grows, leading to a more predictable and uniform structure for the black hole.

### v2

@@ 2 The Model
Reissner-Nordström AdS black holes, which are enveloped by PFDM, carry an electrical charge and are affected by a theoretical type of matter that constitutes the majority of the universe’s mass. These black holes exhibit intriguing thermodynamic behaviors and phase transition characteristics, which can be explored by broadening the phase space of the system. Additionally, some studies have investigated the WCCC in the context of these black holes[114, 115, 116, 117, 118, 119]. This conjecture posits that a black hole’s singularity remains hidden from any external observer. The action under consideration pertains to a gravitational theory that interacts minimally with a gauge field amidst the presence of PFDM[120, 121, 122, 123, 124],
$S=\int d^{4}x\sqrt{-g}\bigg{[}\frac{1}{16\pi G}R+\frac{1}{4}F^{\mu\nu}F_{\mu\nu}-\frac{\Lambda}{8\pi G}+\mathcal{L}_{DM}\bigg{]},$ (2.1)
@@
$\begin{split}R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R+\Lambda g_{\mu\nu}=&-8\pi G(T^{M}_{\mu\nu}+T^{DM}_{\mu\nu})\equiv-8\pi GT_{\mu\nu},\end{split}$ (2.2)
$\begin{split}&F^{\mu\nu}_{;\nu}=0,\\ &F^{\mu\nu;\alpha}+F^{\nu\alpha;\mu}+F^{\alpha\mu;\nu}=0.\end{split}$ (2.3)
Where the second line of Eq.(2.3) is Bianchi identity. The energy-momentum tensor for ordinary matter is symbolized by $T^{M}_{\mu\nu}$, and the energy-momentum tensor for PFDM is represented by $T^{DM}_{\mu\nu}$. These tensors are detailed in[120, 121, 122, 123, 124]. Consequently, the equations can be expressed as follows,
$\begin{split}&T^{\mu}_{\nu}=g^{\mu\nu}T_{\mu\nu},\\ &T^{t}_{t}=-\rho,\hskip 14.22636ptT^{r}_{r}=T^{\theta}_{\theta}=T^{\phi}_{\phi}=P.\end{split}$ (2.4)
Where $\rho$ and $P$ and energy density and pressure, respectively. In the framework of the RN AdS metric influenced by PFDM, it is posited that the components of the energy-momentum tensor satisfy $T^{r}_{r}=T^{\theta}_{\theta}=T^{\phi}_{\phi}=T^{t}_{t}(1-\delta)$, with $\delta$ being a constant. This assumption is substantiated by [109, 120, 122, 123, 124]. The formulation of the RN AdS metric within the PFDM context is delineated as follows[121],
$\begin{split}&ds^{2}=-f(r)dt^{2}+f(r)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),\\ &A_{\mu}=(\frac{Q}{r_{+}},0,0,0),\end{split}$ (2.5)
@@
$f(r)=1-\frac{2MG}{r}+\frac{GQ^{2}}{r^{2}}-\frac{\Lambda}{3}r^{2}+\frac{\gamma}{r}\ln(\frac{r}{|\gamma|}).$ (2.6)
In this analysis, the parameters $M$, $Q$, and $\gamma$ signify the mass, charge and the influence of (PFDM) on the black hole, respectively. Also, $\Lambda=-3/\ell^{2}$ is the cosmological constant with $\ell$ being the AdS radius. In the absence of PFDM ($\gamma=0$), the space-time metric simplifies to that of a standard RN AdS black hole. The parameter $\gamma$ can assume positive values, reflecting different PFDM scenarios. This study contemplates the implications of positive $\gamma$ on the black hole’s characteristics. Also, the extremality bound for charged black holes in AdS without PFDM, $\gamma=0$ reads,
$\begin{split}\frac{|Q|}{\sqrt{G}M}\leq\frac{3|\widetilde{q}|\sqrt{-1+\sqrt{1+2\widetilde{q}^{2}}}}{-1+2\widetilde{q}^{2}+\sqrt{1+2\widetilde{q}^{2}}}\approx\left\{\begin{array}[]{ll}1&\hbox{$|\widetilde{q}|\ll 1$}\\ \frac{3}{(8\widetilde{q}^{2})^{1/4}}&\hbox{$|\widetilde{q}^{2}|\gg 1$,}\end{array}\right.\end{split}$
$\begin{split}\widetilde{q}^{2}:=\frac{6GQ^{2}}{\ell^{2}}.\end{split}$
This nonlinear extremality bound is found by solving $f(r_{\star})\leq 0$ where $r_{\star}$ is the largest solution of f’($r_{\star}$) = 0, where f(r) is given by Eq.(2.6) with $\gamma$ = 0.
@@ 3 RPS thermodynamics
In the RPS formalism, the traditional variables of pressure $P$ and volume $V$ are eschewed due to the altered connotations they acquire within the holographic framework. Instead, two pivotal quantities, $C$ and $\mu$, assume significant roles within the realm of CFT. The central charge $C$ is instrumental in ascertaining the microscopic degrees of freedom inherent in CFT, while its conjugate $\mu$ is regarded as a chemical potential, as referenced in literature[20]. Within the domain of extended phase space thermodynamics, the exploration of black hole thermodynamics necessitates the employment of a dynamic cosmological constant. This adjustment invariably leads to a transformation in the associated dynamical equations. To circumvent this modification, the thermodynamics of constrained phase space is proposed, wherein Newton’s constant is treated as a variable, supplanting the cosmological constant. This substitution preserves the original form of the field equations. Moreover, a correlation is established between the central charge and the cosmological and Newton’s constants, expressed as $C=\frac{\ell^{2}}{G}$, where $\ell^{2}=-\frac{3}{\Lambda}$ [20]. The first law of thermodynamics, as applied to a RN-AdS black hole surrounded by PFDM within the RPS framework, is articulated as follows,
$dM=TdS+\hat{\phi}d\hat{Q}+\mu dC+\hat{\Pi}d\hat{\gamma},$ (3.1)
where, $\hat{\phi}$ and $\hat{Q}$ are the properly re-scaled electric potential electric charge and $\hat{\Pi}$ is the conjugate of the $\hat{\gamma}$. The observation aligns with an Euler-like relation, which is a fundamental principle in thermodynamics and fluid dynamics, often used to describe the relationship between various state functions. In the context of the RPS formalism, such relations are crucial for understanding the thermodynamic properties of black holes and their behavior in different spacetime geometries. Introducing Euler’s relation for RPST can provide a new perspective for studying black hole physics, So,
$M=TS+\hat{\phi}\hat{Q}+\mu C+\hat{\Pi}\hat{\gamma},$ (3.2)
where Eqs.(3.1) and (3.2) are indeed not gauge-invariant at face value. We note that $\hat{\phi}$ represents the difference in electric potential between the black hole horizon and infinity. In the context of our black hole, the parameters $M$, $T$, and $S$ represent the mass, temperature, and entropy, respectively. Moreover, these parameters can generally be characterized by corresponding quantities in the dual CFT,
$\hat{Q}=\frac{Q\ell}{\sqrt{G}}\hskip 28.45274pt\hat{\phi}=\frac{\phi\sqrt{G}}{\ell},\hskip 28.45274pt\hat{\gamma}=\frac{\gamma\ell^{2}}{G},\hskip 28.45274pt\hat{\Pi}=\frac{\Pi G}{\ell^{2}}.$ (3.3)
@@
$C=\frac{\ell^{2}}{G},\qquad\hat{Q}=Q\sqrt{C},\qquad r_{+}=\ell\sqrt{\frac{S}{\pi C}},\qquad\hat{\gamma}=\gamma C.$ (3.4)
Setting $(f(r_{+})=0)$ and utilizing Eqs.(2.6), (3.1), (3.2), and (3.3), we can derive some variables such as M, T, $\hat{\phi}$, $\Pi$, and $\mu$ of the black hole as follows,
$M=\frac{\pi CS+\pi^{2}\hat{Q}^{2}+S^{2}}{2\pi^{3/2}\ell\sqrt{CS}}+\frac{\hat{\gamma}}{2\ell^{2}}\ln\bigg(\sqrt{\frac{SC}{\pi}}\frac{\ell}{\hat{\gamma}}\bigg{missing}),$ (3.5)
$T=\bigg{(}\frac{\partial M}{\partial S}\bigg{)}_{\hat{Q},C,\hat{\gamma}}=\frac{\pi CS-\pi^{2}\hat{Q}^{2}+3S^{2}}{4\pi^{3/2}\ell S\sqrt{CS}}+\frac{\hat{\gamma}}{4\ell^{2}S}$ (3.6)
$\hat{\phi}=\bigg{(}\frac{\partial M}{\partial\hat{Q}}\bigg{)}_{S,C,\hat{\gamma}}=\sqrt{\frac{\pi}{SC}}\frac{\hat{Q}}{\ell},$ (3.7)
$\hat{\Pi}=\bigg{(}\frac{\partial M}{\partial\hat{\gamma}}\bigg{)}_{S,\hat{Q},C}=\frac{1}{2\ell^{2}}\bigg{[}-1+\ln\bigg(\sqrt{\frac{SC}{\pi}}\frac{\ell}{\hat{\gamma}}\bigg{missing})\bigg{]},$ (3.8)
$\mu=\bigg{(}\frac{\partial M}{\partial C}\bigg{)}_{S,\hat{Q},\hat{\gamma}}=\frac{\pi CS-\pi^{2}\hat{Q}^{2}-S^{2}}{4\pi^{3/2}C\ell\sqrt{CS}}+\frac{\hat{\gamma}}{4C\ell^{2}}.$ (3.9)
@@
$\bigg{(}\frac{\partial T}{\partial S}\bigg{)}_{\hat{Q},C,\hat{\gamma}}=0,\qquad\bigg{(}\frac{\partial^{2}T}{\partial S^{2}}\bigg{)}_{\hat{Q},C,\hat{\gamma}}=0.$ (3.10)
Using the above equations for the critical parameters, we have,
$Y:=C\ell^{\frac{4}{3}}\big{(}81(\frac{\hat{\gamma}}{C})^{2}-\ell^{2}+9\ell(\frac{\hat{\gamma}}{C})\sqrt{81(\frac{\hat{\gamma}}{C})^{2}-2\ell^{2}}\big{)}^{\frac{1}{3}},\qquad S_{c}=\frac{\pi(C\ell^{2}+Y)^{2}}{18\ell^{2}Y},$ (3.11)
$\hat{Q}_{c}^{2}=-\frac{\left(C\ell^{2}+Y\right)^{2}\left(C^{2}\ell^{4}-4C\ell^{2}Y+Y^{2}\right)}{324\ell^{4}Y^{2}}+\frac{\sqrt{2}\hat{\gamma}C^{1/2}\left(C\ell^{2}+Y\right)}{9\sqrt{Y}\ell^{2}}.$ (3.12)
@@
$T_{c}=\frac{\left(C\ell^{2}+Y\right)\left(-9\sqrt{2}C^{1/2}\hat{\gamma}\ell^{2}Y^{3/2}+C^{3}\ell^{6}+6C^{2}\ell^{4}Y+6C\ell^{2}Y^{2}+Y^{3}\right)}{3\pi\sqrt{2CY}\ell^{2}\left(C\ell^{2}+Y\right)^{3}}+\frac{9\hat{\gamma}Y}{2\pi\left(C\ell^{2}+Y\right)^{2}},$ (3.13)
$M_{c}=\frac{\sqrt{C}\left(C\ell^{2}+Y\right)\left(2C\ell^{2}+3\hat{\gamma}\sqrt{\frac{2Y}{C}}+2Y\right)}{9\ell^{2}\sqrt{2Y}\left(C\ell^{2}+Y\right)}+\frac{\hat{\gamma}}{2\ell^{2}}\ln\left(\frac{C(C\ell^{2}+Y)}{\hat{\gamma}\sqrt{18CY}}\right).$ (3.14)
@@
$F=M-TS=\frac{\pi CS+3\pi^{2}\hat{Q}^{2}-S^{2}}{4\pi^{3/2}\ell\sqrt{CS}}+\frac{\hat{\gamma}}{4\ell^{2}}\bigg{[}-1+2\ln\bigg(\frac{\ell}{\hat{\gamma}}\sqrt{\frac{SC}{\pi}}\bigg{missing})\bigg{]}.$ (3.15)
Figs.(1) and (2) depict the evolution of free energy as a function of temperature for various free parameter values of the RN AdS black hole, analyzed within the framework of PFDM in RPS. When $\hat{Q}_{c}>\hat{Q}=0.5$, a distinctive swallowtail pattern emerges, indicative of a first-order phase transition. Also, for ($\hat{Q}_{c}=\hat{Q}$ ), we have a second-order phase transition. To enhance clarity, we have separated the figures. The specific free parameters pertinent to each figure are duly annotated. The separation of the figures isolates and emphasizes each parameter set’s unique characteristics, facilitating a more detailed examination of the system’s behavior under varying conditions. It is essential to recognize that the free parameters play a crucial role in shaping the physical properties of the black hole.
Figure 1: With the $\hat{\gamma}=2,C=\ell=1$ and $\hat{Q}_{c}=0.93$ (a) $\hat{Q}_{c}>\hat{Q}=0.5$ (b) $\hat{Q}_{c}=\hat{Q}=0.93$ (c) $\hat{Q}_{c}<\hat{Q}=1.5$
Figure 2: With the $\hat{\gamma}=1.5,C=\ell=1$ and $\hat{Q}_{c}=0.79$ (a) $\hat{Q}_{c}>\hat{Q}=0.5$ (b) $\hat{Q}_{c}=\hat{Q}=0.79$ (c) $\hat{Q}_{c}<\hat{Q}=1.5$
Figure 3: spinodal curve ($T-S$) for charged black hole in presence perfect fluid. We have set $C=\ell=1$ and (a) $\hat{\gamma}=2$ and $\hat{Q}_{c}=0.93$, (b) $\hat{\gamma}=1.5$ and $\hat{Q}_{c}=0.79$
Fig. (3) presents the Hawking temperature as a function of entropy for the AdS (RN) black hole amidst the presence of (PFDM) in the (RPS). We have set the constants $C$ and $\ell$ to 1. For Fig.(LABEL:3a), the parameters are set to $\hat{\gamma}=2$ and $\hat{Q}_{c}$= 0.93 , while for Figure 3b, $\hat{\gamma}=1.5$ and $\hat{Q}_{c}$= 0.79 . Each subplot within the figure reveals certain zero points for varying free parameters within the range $10<S<$ 1. Upon examining the entropy alongside the black hole’s free parameter values, it is observed that for small values of $S$, the structural behavior of the black hole is markedly unique. Conversely, as the value of $S$ increases, the figures tend to converge, indicating a unification in behavior for larger entropy values. This convergence suggests that the influence of the free parameters diminishes as the entropy grows, leading to a more predictable and uniform structure for the black hole.